# Improved Particle Approximation Error for Mean Field Neural Networks

### v2

@@ 5.1 Finite-particle approximation error
We next apply Hoeffding’s inequality to the last term as follows. For any $z\in\mathcal{Z}$, it follows that
$\mathbb{P}_{\mathbf{X}\sim\mu^{\otimes N}_{*}}\left[|h_{\mu_{\mathbf{X}}}(z)-h_{\mu_{*}}(z)|^{2}\geq t\right]\leq 2\exp\left(-\frac{Nt}{2R^{2}}\right).$
By integrating both sides with respect to $t\in[0,\infty)$, we get
$\mathbb{E}_{\mathbf{X}\sim\mu^{\otimes N}_{*}}\left[|h_{\mu_{\mathbf{X}}}(z)-h_{\mu_{*}}(z)|^{2}\right]$ $=\int_{0}^{\infty}\mathbb{P}_{\mathbf{X}\sim\mu^{\otimes N}_{*}}\left[|h_{\mu_{\mathbf{X}}}(z)-h_{\mu_{*}}(z)|^{2}\geq t\right]\mathrm{d}t$
$\leq\int_{0}^{\infty}2\exp\left(-\frac{Nt}{2R^{2}}\right)\mathrm{d}t$
$=\frac{4R^{2}}{N}.$
@@
$N\mathbb{E}_{\mathbf{X}\sim\mu^{\otimes N}_{*}}\left[B_{F}(\mu_{\mathbf{X}},\mu_{*})\right]\leq 2LR^{2}.$
@@
$\frac{1}{N}\mathcal{L}^{(N)}(\mu^{(N)}_{*})-\mathcal{L}(\mu_{*})\leq\frac{2LR^{2}}{N}.$
@@ 5.2 Convergence of Mean-field Langevin dynamics in the discrete-setting
In this section, we showcase auxiliary results used in our theory. The following lemma gives an LSI constant for $\mu^{(N)}_{*}\in\mathcal{P}_{2}(\mathbb{R}^{dN})$.
Lemma 2 (Kook et al., (2024)).
Suppose for any $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$ and $x\in\mathbb{R}^{d}$, $\left|\nabla\frac{\delta F_{0}(\mu)}{\delta\mu}(x)\right|\leq M_{1}$ holds. Then, $\mu^{(N)}_{*}$ satisfies LSI with independent of $N$ given below:
$\bar{\alpha}\geq\frac{\lambda^{\prime}}{8\lambda}\wedge\frac{1}{2}\left(\ \frac{M_{1}}{\lambda^{\prime}}+\sqrt{\frac{\lambda}{\lambda^{\prime}}}\right)^{-2}\left(2+d+\frac{4M_{1}^{2}}{\lambda^{\prime}\lambda}\right)^{-1}\exp\left(-\frac{M_{1}^{2}}{2\lambda^{\prime}\lambda}\right).$
@@
If a probability distribution $\mu\in\mathcal{P}^{2}(\mathbb{R}^{d})$ satisfies the log-Sobolev inequality with constant $\alpha>0$, then $\mu$ satisfies Talagrand’s inequality with the same constant: for any $\mu^{\prime}\in\mathcal{P}_{2}(\mathbb{R}^{d})$

### v3

@@ 5.1 Finite-particle approximation error
The last term is a variance of $h_{\mu_{\mathbf{X}}}(z)~{}(z\in\mathcal{Z})$; hence we simply evaluate it as follows.
$\mathbb{E}_{\mathbf{X}\sim\mu^{\otimes N}_{*}}\left[|h_{\mu_{\mathbf{X}}}(z)-h_{\mu_{*}}(z)|^{2}\right]$ $=\mathbb{E}_{\mathbf{X}\sim\mu^{\otimes N}_{*}}\left[\left|\frac{1}{N}\sum_{i=1}^{N}h(X_{i},z)-\mathbb{E}_{X\sim\mu_{*}}[h(X,z)]\right|^{2}\right]$
$=\mathbb{E}_{\mathbf{X}\sim\mu^{\otimes N}_{*}}\left[\frac{1}{N^{2}}\sum_{i=1}^{N}\left|h(X_{i},z)-\mathbb{E}_{X\sim\mu_{*}}[h(X,z)]\right|^{2}\right]$
$\leq\mathbb{E}_{\mathbf{X}\sim\mu^{\otimes N}_{*}}\left[\frac{1}{N^{2}}\sum_{i=1}^{N}\left|h(X_{i},z)\right|^{2}\right]$
$\leq\frac{R^{2}}{N}.$
@@
$N\mathbb{E}_{\mathbf{X}\sim\mu^{\otimes N}_{*}}\left[B_{F}(\mu_{\mathbf{X}},\mu_{*})\right]\leq\frac{LR^{2}}{2}.$
@@
$\frac{1}{N}\mathcal{L}^{(N)}(\mu^{(N)}_{*})-\mathcal{L}(\mu_{*})\leq\frac{LR^{2}}{2N}.$
@@ 5.2 Convergence of Mean-field Langevin dynamics in the discrete-setting
In this section, we showcase auxiliary results used in our theory.
LSI on $\mu_{*}^{(N)}$ can be verified by using the following two lemmas. Lemma 2 says that the strong log-concave densities satisfy the LSI with a dimension-free constant.
Lemma 2 (Bakry and Émery, (1985)).
Let $\frac{\mathrm{d}\mu(x)}{\mathrm{d}x}\propto\exp(-f(x))$ be a probability density, where $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ is a smooth function. If there exists $c>0$ such that $\nabla^{2}f\succeq cI_{d}$, then $\mu$ satisfies log-Sobolev inequality with constant $c$.
Additionally, the LSI is preserved under bounded perturbation as seen in Lemma 3.
Lemma 3 (Holley and Stroock, (1987)).
Let $\mu$ be a probability distribution on $\mathbb{R}^{d}$ satisfying the log-Sobolev inequality with a constant $\alpha$. For a bounded function $B:\mathbb{R}^{d}\rightarrow\mathbb{R}$, we define a probability distribution $\mu_{B}$ as follows:
$\frac{\mathrm{d}\mu_{B}(x)}{\mathrm{d}x}=\frac{\exp(B(x))}{\mathbb{E}_{X\sim\mu}[\exp(B(X))]}\frac{\mathrm{d}\mu(x)}{\mathrm{d}x}.$
Then, $\mu_{B}$ satisfies the log-Sobolev inequality with a constant $\alpha/\exp(4\|B\|_{\infty})$.
@@
If a probability distribution $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$ satisfies the log-Sobolev inequality with constant $\alpha>0$, then $\mu$ satisfies Talagrand’s inequality with the same constant: for any $\mu^{\prime}\in\mathcal{P}_{2}(\mathbb{R}^{d})$